# Certificate-Sensitive Subset Sum: Realizing Instance Complexity

### v4

@@ 6.2 Refining SETH Lower Bounds
Proof sketch.
If $\epsilon<\delta$, then $O^{*}(2^{\epsilon n})=o^{*}(2^{\delta n})\leq o^{*}(U)$, violating the known SETH lower bound for enumerating $\Sigma(S)$ in the collision-free regime. (See Appendix F for a formal justification of the connection between decision and enumeration hardness in this setting.)
6.3 Revisiting Fine-Grained Reductions
Many standard fine-grained reductions—for example, the classic reduction from CNF-SAT to Multiprocessor Scheduling [6]—produce Subset–Sum instances by mapping boolean literals to powers-of-two. This encoding guarantees that every subset sum is unique, i.e., $U=2^{n}$.
@@ 6.3 Revisiting Fine-Grained Reductions
Take-Away for Fine-Grained Hardness.
Our work shows that fine-grained hardness results for Subset–Sum based on SETH are not truly unconditional. These lower bounds implicitly apply only to instances that are collision-free—i.e., where the number of distinct subset sums $U$ is nearly maximal ($U\approx 2^{n/2}$). Many known reductions, such as those that encode literals via powers-of-two, deliberately create this regime by avoiding arithmetic collisions.
Future reductions must explicitly certify high entropy, or restrict their claims to the collision-free regime. Otherwise, adaptive algorithms like IC-SubsetSum will evade the stated runtime barrier on all structured inputs.
6.4 Guidelines for Future Reductions
To maintain the validity of fine-grained lower bounds, future reductions should satisfy at least one of the following conditions:
  1. G1.
    Collision-resistance certification. Show that the generated instance $(S,t)$ satisfies $|\Sigma(S)|\geq 2^{\delta n}$ for some fixed $\delta>0$.
  2. G2.
    Entropy-aware hardness. Prove hardness not for all of Subset–Sum, but for a carefully defined structural subclass.
  3. G3.
    Non-enumerative goals. If the task is to return a single solution or check feasibility without generating $\Sigma(S)$, ensure that pruning-based methods (like IC-SubsetSum) provably fail on the instance family.
Discussion.
This perspective suggests a simple yet powerful conclusion: any worst-case lower bound for Subset–Sum—or for downstream problems like Knapsack or Scheduling—that fails to account for collision entropy is incomplete. Structural hardness must be made explicit. Otherwise, adaptive algorithms like IC-SubsetSum will continue to evade the theoretical limits by exploiting instance compressibility.
7 Experimental Validation
To validate our theoretical claims, we benchmark our enumerator on synthetic instances where we can precisely control the input structure. We use $n=48$ to allow for complete enumeration runs on a commodity desktop CPU. Our hypothesis is that introducing any form of additive redundancy will cause an exponential collapse in $U=|\Sigma(S)|$, with the runtime of IC-SubsetSum closely tracking this collapse.
We test this by manipulating the input set $S$ along three axes known to induce subset sum collisions: numeric density, duplicate elements, and additive progressions. For each experiment, we run the full enumeration without early termination to measure the total cost of generating the certificate $\Sigma(S)$. Further details and commentary on these experiments are available in Appendix D.
  • Numeric density. Forcing values into a smaller bit-length $w$ creates collisions via the pigeonhole principle.
  • Duplicate elements. Introducing identical elements creates trivial arithmetic dependencies.
  • Additive progressions. Injecting structured sequences creates non-trivial, correlated dependencies.
Table 1: Effect of structure on the size of the distinct sum set $U$. Each entry shows the ratio of observed unique sums to the worst-case maximum $2^{n/2}$ for one half of the set.
Structural Knob Setting $\frac{U}{2^{n/2}}$
Numeric density ($n{=}48$) $w=32\to 24\to 16$ 1.000.840.0271.00$\to$ 0.84$\to$ $0$ 0.84 → 0.027
Duplicates ($n{=}48$) 0240$\to$ 2$\to$ 40 → 2 → 4 duplicates 1.000.570.321.00$\to$ 0.57$\to$ 0.321.00 → 0.57 → 0.32
Additive progressions ($n{=}48$) 1 seq (len 4) / 2 seq (len 4) 0.690.690.69 / 0.470.470.47
Results.
As expected, we observed dramatic variation in the size of $\Sigma(S)$ under different input distributions. Uniform instances often had $U\approx 2^{n/2}$, while arithmetic progressions and structured inputs had $U$ values several orders of magnitude smaller. This confirms that $U$ is not a function of $n$ alone and supports its use as a certificate length parameter.
In all cases, the measured runtime of IC-SubsetSum scaled proportionally with the observed $U$ (for a fixed $n$), which is consistent with the proven $O(U\cdot n^{2})$ complexity bound.
8 Related Work
Subset Sum Algorithms.
The classical Horowitz–Sahni meet-in-the-middle algorithm [8] remains the fastest known worst-case solver for Subset Sum, with $O^{*}(2^{n/2})$ runtime. Despite extensive efforts, no unconditional improvement has been found in general [4, 16]. Our algorithm IC-SubsetSum departs from this tradition by analyzing runtime as a function of the number of distinct subset sums, $U$, rather than the input size $n$.
A recent randomized algorithm by Bringmann, Fischer, and Nakos [4] achieves the first unconditional improvement over Bellman’s dynamic programming algorithm for Subset Sum, solving the target-constrained decision problem in $\widetilde{O}(|S(X,t)|\cdot\sqrt{n})$ time. Their approach focuses on computing reachable sums up to a target $t$, using algebraic and combinatorial techniques. Our algorithm differs in scope: it deterministically enumerates the full set of unique subset sums $\Sigma(S)$, supports canonical constructive witnesses, and adapts its runtime to the instance’s internal structure. While the two approaches solve different variants of Subset Sum, they are complementary in goal and methodology.
Instance Complexity and Adaptive Algorithms.
Instance Complexity was introduced by Orponen, Ker-i, Uwe and Watanabe [10] to capture the inherent difficulty of individual instances. Although the model is powerful in theory, it was widely regarded as non-constructive and unimplementable—a perspective detailed in surveys on the topic (e.g., [5])—offering no path to practical algorithms. This paper directly addresses that long-standing gap. Our result shows, for the first time, how to constructively generate IC certificates online with a runtime that scales with certificate size.
Fine-Grained Complexity.
The fine-grained lower bound framework [15] has produced tight runtime barriers for many problems under assumptions like SETH. Our results add a structural dimension to this picture, showing that hardness results for Subset Sum only apply to collision-free instances. The requirement that reductions certify low entropy has not been previously made explicit. This parallels ideas in compressibility-aware reductions [1], but our formalization via $\Sigma(S)$ and entropy is novel.
Collision Structure in Combinatorics.
The entropy of subset-sum collisions has been studied both in additive combinatorics and in cryptanalytic settings, where low-entropy instances enable significantly faster algorithms. Austrin et al. [2] analyze the structure of subset sums under anti-concentration assumptions, while Howgrave-Graham and Joux [9] apply collision-based meet-in-the-middle techniques to exploit redundancy in integer knapsacks, achieving runtimes below the classical $2^{n/2}$ bound in the randomized setting. These results reinforce the idea that the subset-sum solution space can often be compressed substantially, and that this compressibility—quantified via collision entropy—directly governs algorithmic complexity.
Similarly, classic results in additive combinatorics—such as those related to the Littlewood–Offord problem [14]—bound the number of subset-sums that can collide based on anti-concentration phenomena. Our approach is instance-level and algorithmically constructive: we build data structures that prune repeated sums in real time and link the entropy collapse directly to runtime savings.
9 Future Directions
Closing Algorithmic Gaps for Subset Sum.
Our results open two natural algorithmic questions. First, for our randomized $O(U\cdot n)$ algorithm, can the linear factor in $n$ be eliminated to achieve an $\tilde{O}(U)$ runtime? Second, and more fundamentally, can our deterministic $O(U\cdot n^{2})$ algorithm be improved to match the randomized bound, perhaps by finding a deterministic method for canonical comparison that avoids the $O(n)$ cost?
Structure-Sensitive Runtimes in NP.
Can other NP-complete problems admit certificate-sensitive runtimes? This paper introduces a new structural knob—collision entropy—that complements traditional parameters like solution density. For Partition, Knapsack, and Bin Packing, it may be possible to construct a compressed summary of feasible configurations in time proportional to its size. Formalising such certificates and bounding the corresponding enumeration complexity are exciting open problems.
Derandomization of Local Canonicality.
The randomized $O(U\cdot n)$ algorithm relies on pairwise polynomial hashing to simulate lexicographic comparisons in expected constant time. Is there a deterministic analog that achieves sublinear-time prefix comparison in the local model? Any such result would not only match the conjectured optimal runtime but also resolve a form of local derandomization in a highly practical setting.
Constructive Instance Complexity Beyond Subset Sum.
This paper builds a bridge between Orponen-style instance complexity and efficient enumeration. For Subset–Sum, we show that the certificate $\Sigma(S)$ can be constructed in deterministic $O(U\cdot n^{2})$ time or expected $O(U\cdot n)$ time. Can similar certifiers be constructed online for problems like 3SAT or Clique, and if so, under what entropy or compressibility assumptions? A broader theory of algorithmic IC—one that tracks both information content and construction cost—remains to be written.
Toward a Complexity Class for Certificate-Sensitive Problems.
Our algorithm solves Subset Sum in time polynomial in both the size of its minimal instance-specific certificate, $U=|\Sigma(S)|$, and the input size $n$. This motivates a broader question: are there other NP-complete problems that admit solvers whose runtime is polynomial in the size of an intrinsic structural certificate of the instance? We tentatively propose the name P-IC (Polynomial in Instance Certificate) for such problems. Formalizing this class would require a general definition of “instance certificate”—for example, the set of all satisfying assignments in SAT, or the set of all achievable (value, weight) pairs in Knapsack. Unlike parameterized complexity (e.g., FPT), where the parameter is typically a small, fixed integer supplied with the input, the certificate in our setting is an emergent property of the instance’s internal structure, not known a priori. We believe that defining the boundaries of P-IC, and identifying other problems that fall within it, is a promising direction for future work.
10 Conclusion
Theoretical Advance.
This paper develops a new framework for structure-aware enumeration in NP-complete problems, with Subset Sum as the case study. We introduce a certificate-sensitive complexity measure $U=|\Sigma(S)|$ and show that it aligns with both instance difficulty and practical runtime. Our results offer a unified explanation for why real-world instances are easier, why SETH-based reductions can be structurally brittle, and how instance complexity can be made constructive.
Algorithmic Advance.
IC-SubsetSum achieves a set of performance guarantees that collectively redefine the state of the art:
  • Certificate-sensitive runtime. A deterministic worst-case bound of $O(U\cdot n^{2})$ and an expected randomized bound of $O(U\cdot n)$, where $U$ is discovered online.
  • Worst-case improvement over classic methods, achieving $O^{*}(2^{n/2-\varepsilon})$ for some fixed $\varepsilon>0$.
  • Real-time certificate generation, producing the minimal witness $\Sigma(S)$ with zero redundancy.
Closing the gap between the deterministic and randomized bounds remains a key open challenge, as does the question of whether an $\tilde{O}(U)$ runtime is achievable.
May all your programs be short and quick (Fortnow [5]) and may their runtimes reveal the structure within.
Appendix Overview
Provenance.
This appendix material expands on the enumeration logic, correctness conditions, and structural optimizations introduced across Sections 2–4 and Section 6 of the main paper [11].
We provide the following appendices:
Appendix Overview
This appendix material expands on the enumeration logic, correctness conditions, and structural optimizations introduced in Sections 46 of the main paper.
  • Appendix A provides a detailed description and pseudocode for the core column-wise subset sum enumerator.
  • Appendix B lists the algebraic combination lemmas used by the solver, including correctness checks for sum recombination.
  • Appendix C describes the full double-meet-in-the-middle framework used to exploit structure and symmetry.
  • Appendix D presents additional experimental data, including plots of runtime scaling and collision entropy.
  • Appendix E presents the full Controlled Aliasing framework, including the 4-lane memoization model and implementation-level pseudocode.
  • Appendix F provides a formal justification for the connection between decision and enumeration hardness in high-entropy instances.
Appendix A Column-wise Enumerator: Full Details
Algorithm Description
Column-wise expansion.
Our state space consists of all canonical partial sums $\sigma$ that are still under construction. At each step, we extend $\sigma$ by adding one unused element $a_{i}\notin R$, and we update the prefix to $R\cup\{a_{i}\}$ if the resulting sum $\sigma^{\prime}=\sigma+a_{i}$ is new or establishes a more canonical representation.
This leads to at most $O(U\cdot n)$ total extension attempts from canonical states. The final complexity, as established in Section 4.1, depends on the cost of the canonical comparison performed during collisions.
High-density handling.
To guarantee correctness in high-collision regimes, we use a canonicalization rule based on the lexicographic order of index bitmasks. This ensures that each sum $\sigma$ is associated with exactly one canonical prefix, preventing incorrect pruning due to non-deterministic resolution of duplicated sums.
4-Column Litmus Test.
The real-time nature of our collision detection enables a simple and effective litmus test to estimate an instance’s hardness. By running the enumerator for only a few initial columns (e.g., up to $k=4$)—a process that runs in polynomial time $O(n^{4})$—we can measure the early-stage collision rate among small subsets. A high collision rate strongly predicts that the total number of unique subset sums $U$ will be small, signaling a structurally “easy” instance. Conversely, a low collision rate suggests that $U$ will be large, indicating a harder, unstructured input. This test offers a fast, lightweight method to assess an instance’s intrinsic complexity before engaging in a full exponential-time computation.
A.1 Pseudocode
Algorithm 1 EnumerateUniqueSubsetSums
1:Set $S=\{a_{1},\dots,a_{n}\}$
2:Memo table mapping each unique subset sum to its canonical bitmask.
@@ A.1 Pseudocode
5:for $k=1$ to $n/2$ do
@@
8:   for $i=1$ to $n$ do
9:     if bit $i$ is not set in mask then
10:      new_sum $\leftarrow\texttt{sum}+a_{i}$
11:      new_mask $\leftarrow\texttt{mask}\lor(1\ll(i-1))$
12:      if new_sum $\notin$ Memo then
13:        Memo[new_sum] $\leftarrow$ new_mask
@@
15:      else if IsLexicographicallySmaller(new_mask, Memo[new_sum]then
16:        Memo[new_sum] $\leftarrow$ new_mask
@@
18:  Frontier $\leftarrow$ NextFrontier
19:  if Frontier =$\emptyset$= ∅ then break
20:return Memo
Algorithm 2 IsLexicographicallySmaller
[…]
2:  for $i=1$ to $n$ do
3:   $a_{i}\leftarrow$ bit $i$ of $A$
4:   $b_{i}\leftarrow$ bit $i$ of $B$
5:   if $a_{i}<b_{i}$ then return true
6:   else if $a_{i}>b_{i}$ then return false
7:  return false $\triangleright$ They are equal
A.2 Implementation Invariants
To guarantee correctness and optimal performance, the enumerator must uphold:
  • Deterministic tie-breaking. Among colliding prefixes for a given sum $\sigma$, the lex-minimal one is always retained. This guarantees canonicity across all layers.
  • Efficient storage. Total space is dominated by storing the $U$ canonical prefixes, each represented by an $n$-bitmask. This leads to a space complexity of $O(U\cdot n)$.
  • Duplication elimination. Every sum $\sigma\in\Sigma(S)$ is generated exactly once. This avoids wasted work in high-density regimes and maintains instance-optimality.
Appendix B Compositional Sum Lemmas and Complement Check
@@
then $S\setminus(A\cup B)$ is a witness.
@@
All statements follow from elementary algebra on disjoint subsets and set complements. These identities justify the conditional checks performed by the Check routine in the main solver. Since the enumerator generates all elements of $\Sigma(\ell_{i})$ correctly (§A), all valid witnesses are discoverable by the algorithm.
@@
Appendix C Double-MIM Solver: Detailed Pseudocode and Analysis
C.1 Algorithm Sketch
We assume the set $S=\{a_{1},\dots,a_{n}\}$ is split into left and right halves, $\ell_{0}$ and $\ell_{1}$. While this partition can be arbitrary, its balance can be optimized beforehand via a lightweight, polynomial-time analysis of the instance’s local collision structure (4-Column Litmus Test, Appendix A). We apply the unique-subset-sum enumerator (Appendix A) to $\ell_{0}$ and $\ell_{1}$ to construct their respective certificates, $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$. Then, instead of using a traditional sorted merge, solutions are checked in real time: as each new unique sum is discovered, the algorithm applies the Combination Lemmas and Complement Logic (Appendix B) to determine whether the new sum, or its complement, combines with any previously seen sum from the other split to reach the target $t$.
Algorithm 3 Certificate-Sensitive Subset Sum Solver
1:Multiset $S=\{s_{1},\dots,s_{n}\}$, target $t$
2:Returns a certificate if $t\in\Sigma(S)$; otherwise reports failure
3:Split $S$ into $\ell_{0},\ell_{1}$ alternating by index
4:Input $\leftarrow\{\mathtt{kPerm}(0,\mathbf{0},b)\mid b\in\{0,1\}\}$
[…]
7:for $r=0$ to $n/4$ do
[…]
9:  for all $p\in$ Input do
10:   $L\leftarrow\ell_{p.\texttt{split}}$
[…]
14:     if $p.\texttt{elements}[i]=0$ then
15:      $s^{\prime}\leftarrow p.\texttt{sum}+L[i]$
16:      $e^{\prime}\leftarrow p.\texttt{elements}$ with bit $i$ set to 111
17:      $q\leftarrow\mathtt{kPerm}(s^{\prime},e^{\prime},p.\texttt{split})$
18:      if $s^{\prime}\notin$ Memo then
[…]
21:        NextFrontier $\leftarrow$ NextFrontier $\cup\{q\}$
[…]
24:        Remove $\texttt{Cert}[s^{\prime}]$ from NextFrontier
[…]
27:        NextFrontier $\leftarrow$ NextFrontier $\cup\{q\}$
28:  Input $\leftarrow$ NextFrontier
29:return ”No solution found”
Real-Time Solution Check Procedure (invoked during enumeration):
[…]
@@ C.1 Algorithm Sketch
7:if $s^{\prime}=t$ or $Sum_{b}-s^{\prime}=t$ then
8:  Output: direct solution found
9:  return
10:if $t-s^{\prime}\in$ Memo then
11:  Output: solution across both splits
12:  return
13:if $t-(Sum_{b}-s^{\prime})\in$ Memo then
14:  Output: solution via complement merge
15:  return
C.2 Time and Space Bounds
Let $U_{0}$ and $U_{1}$ be the number of distinct subset sums in the two halves, $\ell_{0}$ and $\ell_{1}$.
Time Complexity.
The runtime of the full solver is governed by the unique-subset-sum enumerator applied to each half of the input. As established in Theorem 1, the deterministic enumerator performs $O(U_{\text{local}}\cdot k^{2})$ work on a set of size $k$ with $U_{\text{local}}$ unique sums. Since our splits are of size $n/2$, each half incurs $O(U_{i}\cdot(n/2)^{2})$ work in the worst case. Although both sets of unique subset sums, $\Sigma(\ell_{0})$ and $\Sigma(\ell_{1})$, are retained in memory to support canonical certificate construction, the runtime is asymptotically dominated by the larger enumerator. This yields a total deterministic complexity of
$O(\max(U_{0},U_{1})\cdot n^{2}).$
The randomized algorithm achieves expected runtime
$O(\max(U_{0},U_{1})\cdot n),$
as shown in Theorem 2.
Space Complexity.
The space is dominated by storing the canonical prefixes for each of the $U_{0}$ and $U_{1}$ states. Each prefix is an $n/2$-bitmask. Therefore, the total space complexity is $O((U_{0}+U_{1})\cdot n)$.
Optimality.
Our randomized algorithm has expected runtime $O(\max(U_{0},U_{1})\cdot n)$, matching the conjectured $\Omega(U\cdot n)$ lower bound from Conjecture 1. This establishes its conditional optimality in expectation. The deterministic algorithm achieves $O(\max(U_{0},U_{1})\cdot n^{2})$ runtime, leaving a polynomial gap relative to the conjectured lower bound. Closing this gap with a deterministic certificate-sensitive algorithm remains an open problem.
Appendix D Experimental Details
We provide additional data and commentary on the experiments presented in Section 7. Our focus is on how structure (e.g., low entropy or repeated values) affects the size of the certificate $\Sigma(S)$, and how this correlates with observed runtime.
D.1 Extremely dense instances ($n=100$)
@@
The data clearly show that the number of distinct subset sums $U$ falls far below the $2^{n/2}$ barrier as $w$ decreases, and the runtime is significantly reduced once $U\ll 2^{n/2}$, consistent with the proven $O(U\cdot n^{2})$ certificate-sensitive bound.
Full experimental results, implementation details, and benchmark data can be found in Section §6 of [11].
D.2 Summary
Across all experiments, we observe that the certificate size $U$ varies by orders of magnitude depending on the input distribution, and the enumerator’s wall-time closely tracks the drop in $U$. This validates the $O(U\cdot n^{2})$ runtime scaling proven in Section 1, as performance for a fixed $n$ is proportional to $U$.
The experiments confirm that IC-SubsetSum is highly adaptive in practice, with performance governed not by the input size alone, but by the structural complexity of the instance itself.
Appendix E Controlled Aliasing: Correctness and Implementation
To ensure correctness under value aliasing, we introduce a 4-lane memoization scheme. Each reachable sum stores up to four canonical bitmasks, corresponding to distinct aliasing configurations. This enables both decision and construction variants of Subset–Sum to be solved without error, even when synthetic collisions are introduced.
Memoization Structure
We define the memoization table as:
$\texttt{Memo}:\mathbb{Z}\rightarrow\left(\{\bot\}\cup\{0,1\}^{n}\right)^{4}$
Each entry maps a sum $\sigma$ to an array of four bitmasks:

### v6

@@ A.2 Implementation Invariants
Appendix B Double-MIM Solver: Detailed Pseudocode and Analysis
B.1 Algorithm Sketch
We split $S=\{a_{1},\ldots,a_{n}\}$ into two halves $\ell_{0}$ and $\ell_{1}$ (sizes within $\pm 1$) and run the unique-subset-sum enumerator on both sides interleaved. Each time a new canonical state is discovered on one half, we immediately invoke a constant-time family of membership tests on the other half (the Check or CheckAliased routines from §C) to detect solutions online. This avoids materializing $\Sigma(S)$ in full.
When Controlled Aliasing is enabled (Section 5.2; Appendix E), each half uses 2-lane memoization keyed by $(\tilde{\sigma},\chi)$ and enforces the count-based canonical normal form (CNF): only masks $b$ with
$c(b):=b[i_{0}]+b[i_{1}]\in\{0,2\}\quad\text{or}\quad\big{(}c(b)=1\ \wedge\ b[i_{1}]=0\big{)}$
are enqueued for expansion; non-canonical states are stored but marked doNotExtend. This yields the $\tfrac{3}{4}$ expansion factor used in the worst-case bound.
B.2 Core Pseudocode (Modular)
We write State(sum, mask, split, doNotExtend) for a per-half enumeration state. Lexicographic comparisons of bitmasks use Algorithm 4 (Appendix A). The solution checks are given in Algorithms 10 and 11 (Appendix C).
Algorithm 6 DoubleMIM-IC-SubsetSum (interleaved, alias-aware)
1:Multiset $S$; target $t$; flag AliasingEnabled;
2:   alias pairs $(x_{0}^{(0)},x_{1}^{(0)})$ at indices $(i_{0}^{(0)},i_{1}^{(0)})$ on $\ell_{0}$, and $(x_{0}^{(1)},x_{1}^{(1)})$ at $(i_{0}^{(1)},i_{1}^{(1)})$ on $\ell_{1}$
3:Returns a witness if $t\in\Sigma(S)$; else reports failure
4:Split $S$ into $\ell_{0},\ell_{1}$ (e.g., alternating indices)
@@ B.2 Core Pseudocode (Modular)
6:Initialize per-half frontiers with the empty state:
7:    Frontier $\leftarrow\{\texttt{State}(0,\mathbf{0},0,\texttt{false}),\ \texttt{State}(0,\mathbf{0},1,\texttt{false})\}$
8:Insert the empty key on both halves:
9:    $(0,\mathbf{0},0);  InsertOrImprove$$(0,\mathbf{0},1)$
10:for $r=0$ to $\lfloor n/2\rfloor$ do $\triangleright$ interleave by “column”/Hamming weight
11:  NextFrontier $\leftarrow\emptyset$
12:  for all $p\in$ Frontier do
13:   if $p.\texttt{doNotExtend}$ then continue
14:   $b\leftarrow p.\texttt{split}$;  $L\leftarrow\ell_{b}$;  $k\leftarrow|L|$
15:   for $i=0$ to k1k-1 do
16:     if $p.\texttt{mask}[i]=0$ then
17:      $s^{\prime}\leftarrow p.\texttt{sum}+L[i]$;  $e^{\prime}\leftarrow p.\texttt{mask}$ with bit $i\leftarrow 1$
[…]
19:  Frontier $\leftarrow$ NextFrontier
20:  if Frontier == $\emptyset$ then break
21:return “No solution found”
[…]
1:Half index $b\in\{0,1\}$; candidate sum $s^{\prime}$ and mask $e^{\prime}$
2:/* Compute per-half key (aliased or not) */
@@
5:if wasNewwasNew then
6:  $q\leftarrow\texttt{State}(s^{\prime},e^{\prime},b,\texttt{false})$
[…]
8:   $q.\texttt{doNotExtend}\leftarrow\texttt{true}$
9:  NextFrontier $\leftarrow$ NextFrontier $\cup\{q\}$
10:  if AliasingEnabled then CheckAliased(q)(q) else Check(q)(q)
[…]
1:Half index $b$; real sum $s^{\prime}$; mask $e^{\prime}$
2:if AliasingEnabled then
3:  $\chi\leftarrow e^{\prime}[\,i_{1}^{(b)}\,]$
4:  $\tilde{s}\leftarrow s^{\prime}-\chi\cdot\big{(}x_{1}^{(b)}-x_{0}^{(b)}\big{)}$
5:  return $(\tilde{s},\chi)$
6:else
7:  return $s^{\prime}$
[…]
1:Key = $s^{\prime}$ (no alias) or $(\tilde{s},\chi)$ (alias); mask $e^{\prime}$; half $b$
[…]
3:if $\texttt{key}\notin\texttt{Memo}$ then
@@
6:  return true $\triangleright$ new canonical representative
7:else if IsLexicographicallySmaller($e^{\prime}$, Memo[key]) then
8:  /* suppress old representative so it never extends further */
9:  if $\texttt{Rep}[\texttt{key}]\neq\texttt{nil}$ then   $\texttt{Rep}[\texttt{key}].\texttt{doNotExtend}\leftarrow\texttt{true}$
@@
12:  return true $\triangleright$ canonical representative updated
13:else
14:  return false
Notes on Modularity.
  • Enumeration Logic. DoubleMIM-IC-SubsetSum drives interleaved enumeration; ProcessExtension encapsulates “generate, deduplicate, (optionally) gate by CNF, then check”.
  • Deduplication. InsertOrImprove implements the lex-minimal canonicalization with suppression-on-overwrite, ensuring each canonical key is expanded at most once.
  • Aliasing Hooks. KeyOf provides a single switch for aliased vs. true sums; the CNF gate is the call to Expandable()($\cdot$) (Appendix E).
  • Solution Checks. The real-time checks defer to Algorithms 10 and $1$.
B.3 Time and Space Bounds
Let $U_{i}:=|\Sigma(\ell_{i})|$ and $\widehat{U}:=\max(U_{0},U_{1})$.
Without Aliasing.
By Theorems 1 and 2, enumerating each half in isolation costs $O(U_{i}\cdot(n/2)^{2})$ deterministically, and $O(U_{i}\cdot(n/2))$ in expectation for the randomized variant. In the interleaved solver, every cross-half probe and canonical comparison can be injectively charged to a unique successor attempt on the dominant side (the half with $U_{i}=\widehat{U}$) due to suppression-on-overwrite. Hence
$T_{\mathrm{det}}(n)=O\!\big{(}\widehat{U}\cdot n^{2}\big{)},\qquad\mathbb{E}\big{[}T_{\mathrm{rand}}(n)\big{]}=O\!\big{(}\widehat{U}\cdot n\big{)},$
which is asymptotically equivalent to the sum form $O\!\big{(}(U_{0}{+}U_{1})\cdot n^{2}\big{)}$ and $O\!\big{(}(U_{0}{+}U_{1})\cdot n\big{)}$ stated in Theorem 4. Space is $O\!\big{(}(U_{0}{+}U_{1})\cdot n\big{)}$ for storing canonical bitmasks.
With Controlled Aliasing (CNF).
Let $E_{i}$ be the number of expanded states on half $i$ under CNF. By Lemma 9, $E_{i}\leq\tfrac{3}{4}\cdot 2^{|\ell_{i}|}=\tfrac{3}{4}\cdot 2^{n/2}$ in the worst case. Substituting $E_{i}$ for $U_{i}$ in the same accounting yields
$T_{\mathrm{det}}(n)=O\!\big{(}\max(E_{0},E_{1})\cdot n^{2}\big{)}=O^{*}\!\left(2^{\,n/2-\log_{2}(4/3)}\right),$
$\qquad\mathbb{E}\big{[}T_{\mathrm{rand}}(n)\big{]}=O\!\big{(}\max(E_{0},E_{1})\cdot n\big{)}=O^{*}\!\left(2^{\,n/2-\log_{2}(4/3)}\right),$
matching Theorem 6. Space remains $O^{*}(2^{n/2})$, since each aliased key stores up to two canonical masks (one per lane $\chi\in\{0,1\}$).
Optimality (Conditional).
Under Conjecture 1, the randomized solver is optimal up to constants in the local deduplicating model, achieving $\Theta(U\cdot n)$ expected time (Theorem 3). The deterministic $O(U\cdot n^{2})$ bound leaves a polynomial gap, whose removal would require a deterministic sublinear-time canonicality test per extension.
Summary. The interleaved double–meet-in-the-middle framework, combined with (i) lex-minimal canonicalization with suppression-on-overwrite and (ii) the CNF expansion gate under Controlled Aliasing, yields (a) certificate-sensitive runtimes scaling with $U_{0},U_{1}$ and (b) a strict worst-case improvement below $2^{n/2}$ on all inputs.
Appendix C Compositional Sum Lemmas and Complement Check
@@
then $S\setminus(A\cup B)$ is a witness (i.e., sums to $t$). Equivalently,
$\sigma(\ell_{0}\setminus A)+\sigma(\ell_{1}\setminus B)=\sigma(S)-t\ \Longleftrightarrow\ \sigma(A)+\sigma(B)=t.$
@@
All statements follow from elementary algebra on disjoint subsets and set complements (e.g., $\sigma(\ell_{i}\setminus A)=\sigma(\ell_{i})-\sigma(A)$ and $\ell_{0}$ is disjoint from $\ell_{1}$). These identities justify the conditional checks performed by the Check routine in the main solver. Since the enumerator generates all elements of $\Sigma(\ell_{i})$ correctly (§A), all valid witnesses are discoverable by the algorithm.
@@
Real-Time Solution Check Procedures (invoked during enumeration).
Algorithm 10 Check(q)(q) — standard (no aliasing)
1:$q$ — a $k$-permutation state
@@
8:if $s_{\mathrm{real}}=t$ or $\mathrm{Sum}_{b}-s_{\mathrm{real}}=t$ then
9:  Output: direct solution found
10:  return Cross-half patterns:
11:(D)       check $t-s_{\mathrm{real}}\in$ Memo                         $\triangleright$ $\sigma(A)+\sigma(B)=t$
12:if $t-s_{\mathrm{real}}\in$ Memo then
13:  Output: solution across splits; return
14:(RC)   check $\mathrm{Sum}_{b^{\prime}}-(t-s_{\mathrm{real}})\in$ Memo $\triangleright$ $\sigma(A)+\sigma(\ell_{b^{\prime}}\!\setminus\!B)=t$
15:if $\mathrm{Sum}_{b^{\prime}}-(t-s_{\mathrm{real}})\in$ Memo then
16:  Output: mixed (right-complement); return
[…]
18:(LC)   check $t-s_{\mathrm{comp}}\in$ Memo                        $\triangleright$ $\sigma(\ell_{b}\!\setminus\!A)+\sigma(B)=t$
19:if $t-s_{\mathrm{comp}}\in$ Memo then
20:  Output: mixed (left-complement); return
21:(DC)   check $\mathrm{Sum}_{b^{\prime}}-(t-s_{\mathrm{comp}})\in$ Memo $\triangleright$ $\sigma(\ell_{b}\!\setminus\!A)+\sigma(\ell_{b^{\prime}}\!\setminus\!B)=t$
22:if $\mathrm{Sum}_{b^{\prime}}-(t-s_{\mathrm{comp}})\in$ Memo then
23:  Output: double complement; return
Algorithm 11 CheckAliased(q)(q) — 2-lane $(\tilde{\sigma},\chi)$
1:$q$ — a $k$-permutation state; alias pairs $(x_{0}^{(0)},x_{1}^{(0)})$ on $\ell_{0}$ and $(x_{0}^{(1)},x_{1}^{(1)})$ on $\ell_{1}$ with indices $(i_{0}^{(0)},i_{1}^{(0)})$, $(i_{0}^{(1)},i_{1}^{(1)})$
2:Verifies whether $q$ completes a valid solution under Controlled Aliasing
3:Let $t$ be the target